OpenStudy (anonymous):

On my paper it says to use Pythagorean Theorem to find the distance,to the nearest tenth,between F(9,5) and G(-2,2). how do i do this?

OpenStudy (anonymous):

from \(-2\) to \(9\) is \(11\) units to the right as \(9-(-2)=11\) from \(2\) to \(5\) is \(3\) units up as \(5-2=3\) then by pythagoras, the diagonal is \(\sqrt{11^2+3^2}\)

OpenStudy (anonymous):



OpenStudy (anonymous):

i got 14.6 units?

OpenStudy (anonymous):

i got \(11.4\)